# Edge Computing based Human-Robot Cognitive Fusion: A Medical Case Study in the Autism Spectrum Disorder Therapy

### v1

@@ Abstract
In recent years, edge computing has served as a paradigm that enables many future technologies like AI, Robotics, IoT, and high-speed wireless sensor networks (like 5G) by connecting cloud computing facilities and services to the end users. Especially in medical and healthcare applications, it provides remote patient monitoring and increases voluminous multimedia. From the robotics angle, robot-assisted therapy (RAT) is an active-assistive robotic technology in rehabilitation robotics, attracting many researchers to study and benefit people with disability like autism spectrum disorder (ASD) children. However, the main challenge of RAT is that the model capable of detecting the affective states of ASD people exists and can recall individual preferences. Moreover, involving expert diagnosis and recommendations to guide robots in updating the therapy approach to adapt to different statuses and scenarios is a crucial part of the ASD therapy process. This paper proposes the architecture of edge cognitive computing by combining human experts and assisted robots collaborating in the same framework to help ASD patients with long-term support. By integrating the real-time computing and analysis of a new cognitive robotic model for ASD therapy, the proposed architecture can achieve a seamless remote diagnosis, round-the-clock symptom monitoring, emergency warning, therapy alteration, and advanced assistance.
Introduction
Edge Computing offers the full computation or part of the computation that can process the data at the edge network, which enables low latency, faster response, and more comprehensive data analysis (Khan et al. 2019). The connected devices can provide services in AI, robotics, autonomous driving, smart cities, healthcare, medical diagnosis, smart grids, multimedia, and security through the edge network. As one of the significant components of the smart city, the field of smart healthcare emerges from the need to improve the management of the healthcare sector, better utilize its resources, and reduce its cost while maintaining or even enhancing its quality level (Oueida et al. 2018). Traditional smart healthcare systems can be divided into three layers: the collection layer (gathering the sensing data from patients), the transmission layer (sending the data to the base station through the intelligent terminal), and the analysis layer (analyzing the data in the cloud data center), which lack of real-time monitor, emergency service, and comprehensive disease analysis, high communication latency, inflexible network resource deployment, etc (Chen et al. 2018). Although the recent 5G network can support edge computing–based healthcare systems, several challenges still hinder its application and benefits to the entire human community, such as large-scale healthcare, big data management, and patient information privacy (Hartmann, Hashmi, and Imran 2022).
@@ Introduction
Especially when integrating AI and robotics technology, future medical and healthcare applications will involve vast amounts of real-time clinical data computation and analysis. Moreover, combining human experts’ feedback based on real-time results from intelligent terminals is also critical for the future of medical systems and smart healthcare, which can provide professional recommendations, personalized services, and corresponding accurate medical treatment measures. Taking robot-assisted therapy (RAT) for autism spectrum disorder (ASD) [^1] ASD is a developmental disability caused by differences in the brain, which affects the patients’ normal interactions, such as learning, moving, communicating, and paying attention. as an example, it is essential to enable RAT to adapt to individual unique and changing needs. (Clabaugh et al. 2019) formalized personalization as a hierarchical human-robot learning framework consisting of five controllers mediated by a meta-controller that utilized reinforcement learning to personalize instruction challenge levels and robot feedback based on each user’s unique learning patterns. To some extent, although it achieves long-term in-home deployments with children with ASD, it is hard to clarify the role of human experts in the therapy section and further improve their social skills in the real world. Therefore, the quality of the machine-to-machine communication and the human-robot interaction data fusion is the pre-condition for providing efficient and effective medical service, which can improve patient care experience and increase flexibility and adaptability from a business perspective (Wan, Gu, and Ni 2020).
@@
From the cognitive computing perspective, (Chen et al. 2018) introduces the Edge-Cognitive-Computing-based (ECC-based) smart healthcare system, which can monitor and analyze users’ physical health using cognitive computing and solve the problems of inflexible network resource deployment. (Yvanoff-Frenchin et al. 2020) develops a multi-language robot interface based on edge computing, helping evaluate seniors’ mental health by interacting through questions. However, from the edge intelligent devices (like robots) angle (Groshev et al. 2023), there are still several open questions in RAT for healthcare, such as what the best roles for robots are in therapy, how to develop a general approach to integrate robots into interventions adapting various patients’ needs and recognize their status, and who among individuals with the specific symptoms are best suited for this approach, especially in RAT for ASD (Diehl et al. 2012).
@@
From the system design perspective, edging computing involves complex and heterogeneous architecture, which makes it hard to build a general framework for every edging computing application ultimately (Krishnasamy, Varrette, and Mucciardi 2020). Particularly, smart healthcare systems are dynamic, flexible, and complex systems with unpredictable behaviors and need to organize resources and personalize diverse services efficiently (Oueida et al. 2018).
@@
In order to address those challenges, this paper proposes the architecture of the edge computing by combining human experts and assisted robots collaborating in the same framework to help ASD patients with long-term support. By integrating the real-time computing and analysis of a new cognitive robotic model for ASD therapy based on cognitive-behavior therapy (CBT) (Beck 2020) with the proposed architecture, it can achieve a seamless remote diagnosis, round-the-clock symptom monitoring, emergency warning, therapy alteration, and advanced assistance. Fig. 1 outlines the proposed architecture of edge cognitive computing and we list the main contributions of this research as follow:
@@
This section briefly reviews the edge computing and cognitive robotics, and provide a brief background to the cognitive-behavior therapy (CBT). When describing a specific method, we use the notations and relative definitions from the corresponding papers.
Edge Computing
In recent years, edge computing has served as a critical enabler for lots of future technologies like 5G, the Internet of Things (IoT), augmented reality, and cooperative multi-agent/robot systems (such as vehicle-to-vehicle communications by connecting cloud computing facilities and services to end users), which brings the service and utilities of cloud computing closer to the end user and is characterized by fast processing and quick application response time (Khan et al. 2019). Specifically, in order to address the high latency between cloud servers and end-user devices in different scenarios, edge computing tackles them by bringing the processing to the edge of the network (like 5G) and building the corresponding edge computing models, such as Cloudlets (Shaukat et al. 2016), Fog computing (Bao et al. 2017), and Mobile Edge computing (Ahmed and Rehmani 2017), solving cloud computing issues.
Furthermore, with widely implemented deep learning (DL) models, including computer vision and natural language processing, in current end devices (like smart phone, mobile robot, and IoT devices), analyzing their generated data in real-time is essential for system performance and user experience. Therefore, edge computing, where a fine mesh of compute nodes are placed close to end devices, is a viable way to meet the high computation and low-latency requirements of DL on edge devices and also provides additional benefits in terms of privacy, bandwidth efficiency, and scalability (Chen and Ran 2019). As the representative technique of artificial intelligence (AI), integrating DL into edge computing frameworks can build an intelligent edge for dynamic, adaptive edge maintenance and management, improving the edge intelligent agents’ capabilities and extending their range of applications essentially (Wang et al. 2020).
Cognitive Robotics
Figure 2: The illustration of the model for a generic intrinsically motivated agent.
@@
Cognitive robotics studies the mechanisms, architectures, and constraints that allow lifelong and open-ended improvement of perceptual, reasoning, planning, social, knowledge acquisition, and decision-making skills in embodied machines (Merrick 2017). Since cognitive modeling uses symbolic coding schemes depicting the world, perception, action, and symbolic representation become the core issues in cognitive robotics (Yang et al. 2019; Yang and Parasuraman 2020a, 2021). Related disciplines are not just artificial intelligence and robotics but also neuroscience, cognitive science, developmental psychology, sociology, and so on (Asada et al. 2009). Especially building a value system to mimic the “brain” of an AI agent mapping behavioral responses for sensed external phenomena is the core component of cognitive robotics, which is also an emerging and specialized sub-field in neurorobotics, robotics, and artificial cognitive systems research (Yang and Liu 2023). Fig. 2 presents a model for a generic intrinsically motivated agent in cognitive robotics (Merrick 2013).
@@
Here, the value measures an agent’s effort to expend to obtain a reward or avoid punishment (Yang and Parasuraman 2020b). It is not hard-wired for an AI agent, even a biological entity, and the specific value system achieved through experience reflects an agent’s subjective evaluation of the sensory space (Yang and Parasuraman 2023a). And the value mechanisms usually have been defined as the expected values, particularly in uncertain environments. Moreover, the innate value reflects an agent’s subjective evaluation of the sensory space, but the acquired value is shaped through experience during its development.
@@
According to the development of cognitive robotics, the existing value systems can mainly be classified into three categories. Neuroanatomical Systems discuss the explainable biologically inspired value systems design from neuroanatomy and physiology perspectives (Sporns, Almássy, and Edelman 2000); Neural Networks Systems build more abstract models through mathematical approaches to mimic the agent’s value systems; Motivational Systems consider the model that agents interact with environments to satisfy their innate values, and the typical mechanism is reinforcement learning (RL) (Yang and Parasuraman 2023b; Yang 2023).
Cognitive-Behavior Therapy
@@
Cognitive behavioral therapy (CBT) is a type of psychotherapeutic treatment that helps people learn how to identify and change destructive or disturbing thought patterns that negatively influence their behavior and emotions (Hofmann et al. 2012). It is based on the cognitive model (Ellis 1962), which hypothesizes that people’s emotions, behaviors, and physiology are influenced by their perception of events. In all forms of CBT that are derived from Beck’s model (Beck 1964), treatment is based on a cognitive formulation, beliefs, and behavioral strategies that characterize a specific disorder (Alford, Beck, and Jones Jr 1997). Fig. 3 illustrates the relationship of behavior to automatic thoughts with the hierarchy of cognition (Beck 2020).
Figure 3: Illustration of the relationship of behavior to automatic thoughts with the hierarchy of cognition.
@@
CBT involves a wide range of strategies to help people overcome negative patterns, such as identifying negative thoughts, practicing new social skills, SMART goal-setting, stress problem-solving, and self-monitoring (Beck 2020). It can be used as a short-term treatment to help individuals learn to focus on present thoughts and beliefs (symptoms like addiction, anger issues, anxiety, depression, etc.) and improve mental health conditions (like chronic pain, insomnia, stress management, etc.) (Hofmann et al. 2012).
Approach Overview
@@
Cognitive Computing Architecture
@@
Through receiving the analysis results from the cognitive data server, this server will learn resources about edge cloud computing, network communication, and medicine. Then, it optimizes scheduling strategies and distributes corresponding resources in real-time. Moreover, the server sends the integrated resource data back to the cognitive data server, updating its database. Specifically, the server can achieve resource optimization and energy saving by implementing the computing unload, handover strategy, caching and delivery, and the corresponding intelligent algorithms (Chen et al. 2018), which can meet various heterogeneous application requirements.
Edge Layer
Instead of constantly moving data to the cloud for computing operations, which accounts for the energy costs, in the edge layer, data can be mined and processed on edge devices and servers closer to the user (Bhargava and Ivanov 2017). Moreover, compared with traditional edge devices, we define a novel edge intelligent robot for medical applications and smart health care, which not only can collect and process data from users and the environment but also serve as the medical equipment or tool to cure patients.
@@
Cognitive Models of the Edge Intelligent Robot
@@
Humor Styles for Cognitive Distortions
The type of humor reflects the cognitive development of people and the level of their social skills (Bernet 1993). Humor styles are potential mediators of the association between cognitive and interpersonal vulnerability factors and psychological dysfunction, distress, or poor interpersonal functioning (Rnic, Dozois, and Martin 2016). To some extent, by improving their sense of humor, ASD patients can gradually enhance their communication skills and enter the mutual socialization process with others (Southam 2005).
Based on what McGhee (McGhee and Pistolesi 1979) described, this research considers four stages of humor development in the whole CBT process. Using the assisted robot to design different funny scenarios and jokes treats ASD patients, especially the children, in their corresponding periods. Tab. 1 illustrates the four stages of human development through RAT in our research.
Table 1: Four Stages of Humor Development through RAT
@@
Behavior Tree based Humor Styles Representation
According to the above discussion, we design four specific application cases representing them as behavior trees (BT) (Colledanchise and Ögren 2018) to improve patients’ experience and enhance the effect of the RAT in different stages of ASD.
@@
Patients with less social skills enjoy interactions with familiar objects or games by creating something new or different out of them. Therefore, at the entry level, we design three scenarios (playing ball, chasing, and spinning) to let the assisted robot interact with patients, helping them develop interests and willingness to communicate and make new friends. Fig. 4 shows the three scenarios as the BT.
[…]
Figure 4: Entry Level: Assisted Robot’s Behavior Trees
@@
As patients develop interests and are not resistant to communicating with new partners, they become willing to use language and gestures for fun and to engage others. There may be some overlap between Entry Level and Basic Level in humor development, but the distinguishing feature of the Basic Level is that a ”verbal statement alone creates the incongruity and leads to laughter” (McGhee and Pistolesi 1979).
At this point, humor becomes a significant part of social skill development. In order to help patients gain positive emotional reactions from their partners, we create a simple pretend game  [^2] It is a loosely structured form of play that generally includes role-play, object substitution, and nonliteral behavior (Fein 1981). – Aladdin and the Magic Lamp [^3] Here, the assisted robot plays the Magic Lamp and the patient plays the Aladdin. – to let them play with the robot, adding social reinforcement to the fun. Fig. 5 illustrates an example of the Assisted Robot’s Behavior Trees of “Aladdin and the Magic Lamp”.
ptRAT: Basic Level${}~{}:What~{}do~{}you~{}want?~{}~{}$pt$:[The~{}Star!]$$:Are~{}you~{}afraid~{}to~{}fly?$pt$:[Yes!]$$:Better~{}in~{}earth!$pt$:[No!]$$:Let^{\prime}s~{}go!$pt$:[Others]$$:Sorry,$$Out~{}of~{}store!$
Figure 5: Entry Level: an Example of the Assisted Robot’s Behavior Trees of “Aladdin and the Magic Lamp”.
@@
When patients master the above two levels of humor, their cognition will develop to Piaget’s Preoperational Stage (Papalia, Olds, and Feldman 2007). Humor can be regarded as the intellectual play expressed through language (Freud 1960), and the production and appreciation of humor also change.
In this intellectual linguistic play, the listener needs to understand the double meanings of words to find it humorous (Southam 2005). At this stage, making others laugh becomes the reward of social approval, which can drive patients to develop this social language skill. Knock-knock jokes and other ready-made jokes are popular at this level. Considering knock-knock jokes and other ready-made jokes are popular, we design the scenarios of knock-knock jokes to support patients interacting with the robot at this level (Fig. 6).
[…]
Figure 6: Middle Level: an Example of the Assisted Robot’s Behavior Trees of the Knock knock joke.
@@
In the cognitive stage of concrete operations (Papalia, Olds, and Feldman 2007), patients will improve mental operations to analyze multiple aspects of a situation and perform tasks at a higher level than they could in the preoperational stage. They need to understand the meanings of puns and other forms of more abstract humor and can use inductive and deductive reasoning and reversibility in thinking about the beginning, middle, and end points of a funny story or joke (Southam 2005). Teasing and sarcastic jokes are commonly used at this level. Fig. 7 illustrates two examples of assisted robot’s BT of sarcastic jokes.
ptRAT: Advanced Level${}~{}:Do~{}you~{}like~{}me?~{}~{}$pt$:[Yes!]$$:You^{\prime}re~{}a~{}wise$$person.$pt$:[No!]$$:Not~{}everyone$$has~{}good~{}taste!$
ptRAT: Advanced Level${}~{}:Do~{}you~{}find~{}me~{}offensive?~{}~{}$pt$:[Yes!]$$:I~{}suggest~{}you$$quit~{}finding~{}me.$pt$:[No!]$$:Enjoy~{}it!$
Figure 7: Advanced Level: Examples of Assisted Robot’s Behavior Trees of the Sarcastic Jokes.
Conclusion and Future Works

### v2

@@ Abstract
In recent years, edge computing has served as a paradigm that enables many future technologies like AI, Robotics, IoT, and high-speed wireless sensor networks (like 5G) by connecting cloud computing facilities and services to the end users. Especially in medical and healthcare applications, it provides remote patient monitoring and increases voluminous multimedia. From the robotics angle, robot-assisted therapy (RAT) is an active-assistive robotic technology in rehabilitation robotics, attracting researchers to study and benefit people with disability like autism spectrum disorder (ASD) children. However, the main challenge of RAT is that the model capable of detecting the affective states of ASD people exists and can recall individual preferences. Moreover, involving expert diagnosis and recommendations to guide robots in updating the therapy approach to adapt to different statuses and scenarios is a crucial part of the ASD therapy process. This paper proposes the architecture of edge cognitive computing by combining human experts and assisted robots collaborating in the same framework to achieve a seamless remote diagnosis, round-the-clock symptom monitoring, emergency warning, therapy alteration, and advanced assistance.
I Introduction
Edge Computing offers the full computation or part of the computation that can process the data at the edge network, which enables low latency, faster response, and more comprehensive data analysis [1]. The connected devices can provide services in AI, robotics, autonomous driving, smart cities, healthcare, medical diagnosis, smart grids, multimedia, and security through the edge network. As one of the significant components of the smart city, the field of smart healthcare emerges from the need to improve the management of the healthcare sector, better utilize its resources, and reduce its cost while maintaining or even enhancing its quality level [2]. Traditional smart healthcare systems can be divided into three layers: the collection layer (gathering the sensing data from patients), the transmission layer (sending the data to the base station through the intelligent terminal), and the analysis layer (analyzing the data in the cloud data center), which lack of real-time monitor, emergency service, and comprehensive disease analysis, high communication latency, inflexible network resource deployment, etc [3]. Although the recent 5G network can support edge computing–based healthcare systems, several challenges still hinder its application and benefits to the entire human community, such as large-scale healthcare, big data management, and patient information privacy [4].
@@ I Introduction
Especially when integrating AI and robotics technology, future medical and healthcare applications will involve vast amounts of real-time clinical data computation and analysis. Moreover, combining human experts’ feedback based on real-time results from intelligent terminals is also critical for the future of medical systems and smart healthcare, which can provide professional recommendations, personalized services, and corresponding accurate medical treatment measures. Taking robot-assisted therapy (RAT) for autism spectrum disorder (ASD) [^1] ASD is a developmental disability caused by differences in the brain, which affects the patients’ normal interactions, such as learning, moving, communicating, and paying attention. as an example, it is essential to enable RAT to adapt to individual unique and changing needs. [5] formalized personalization as a hierarchical human-robot learning framework consisting of five controllers mediated by a meta-controller that utilized reinforcement learning to personalize instruction challenge levels and robot feedback based on each user’s unique learning patterns. To some extent, although it achieves long-term in-home deployments with children with ASD, it is hard to clarify the role of human experts in the therapy section and further improve their social skills in the real world. Therefore, the quality of the machine-to-machine communication and the human-robot interaction data fusion is the pre-condition for providing efficient and effective medical service, which can improve patient care experience and increase flexibility and adaptability from a business perspective [6].
@@
From the cognitive computing perspective, [3] introduces the Edge-Cognitive-Computing-based (ECC-based) smart healthcare system, which can monitor and analyze users’ physical health using cognitive computing and solve the problems of inflexible network resource deployment. [7] develops a multi-language robot interface based on edge computing, helping evaluate seniors’ mental health by interacting through questions. However, from the edge intelligent devices (like robots) angle [8], there are still several open questions in RAT for healthcare, such as what the best roles for robots are in therapy, how to develop a general approach to integrate robots into interventions adapting various patients’ needs and recognize their status, and who among individuals with the specific symptoms are best suited for this approach, especially in RAT for ASD [9].
@@
From the system design perspective, edging computing involves complex and heterogeneous architecture, which makes it hard to build a general framework for every edging computing application ultimately [10]. Particularly, smart healthcare systems are dynamic, flexible, and complex systems with unpredictable behaviors and need to organize resources and personalize diverse services efficiently [2].
@@
In order to address those challenges, this paper proposes the architecture of the edge computing by combining human experts and assisted robots collaborating in the same framework to help ASD patients with long-term support. By integrating the real-time computing and analysis of a new cognitive robotic model for ASD therapy based on cognitive-behavior therapy (CBT) [11] with the proposed architecture, it can achieve a seamless remote diagnosis, round-the-clock symptom monitoring, emergency warning, therapy alteration, and advanced assistance. Fig. 1 outlines the proposed architecture of edge cognitive computing and we list the main contributions of this research as follow:
@@ II Background and Preliminaries
This section briefly reviews cognitive robotics and provide a brief background to the cognitive-behavior therapy (CBT). When describing a specific method, we use the notations and relative definitions from the corresponding papers.
II-A Cognitive Robotics
@@
Cognitive robotics studies the mechanisms, architectures, and constraints that allow lifelong and open-ended improvement of perceptual, reasoning, planning, social, knowledge acquisition, and decision-making skills in embodied machines [12]. Since cognitive modeling uses symbolic coding schemes depicting the world, perception, action, and symbolic representation become the core issues in cognitive robotics [13, 14, 15]. Related disciplines are not just artificial intelligence and robotics but also neuroscience, cognitive science, developmental psychology, sociology, and so on [16]. Especially building a value system to mimic the “brain” of an AI agent mapping behavioral responses for sensed external phenomena is the core component of cognitive robotics, which is also an emerging and specialized sub-field in neurorobotics, robotics, and artificial cognitive systems research [17].
@@
Here, the value measures an agent’s effort to expend to obtain a reward or avoid punishment [18]. It is not hard-wired for an AI agent, even a biological entity, and the specific value system achieved through experience reflects an agent’s subjective evaluation of the sensory space [19]. And the value mechanisms usually have been defined as the expected values, particularly in uncertain environments. Moreover, the innate value reflects an agent’s subjective evaluation of the sensory space, but the acquired value is shaped through experience during its development.
@@
According to the development of cognitive robotics, the existing value systems can mainly be classified into three categories. Neuroanatomical Systems discuss the explainable biologically inspired value systems design from neuroanatomy and physiology perspectives [20]; Neural Networks Systems build more abstract models through mathematical approaches to mimic the agent’s value systems; Motivational Systems consider the model that agents interact with environments to satisfy their innate values, and the typical mechanism is reinforcement learning (RL) [21, 22, 23].
II-B Cognitive-Behavior Therapy
@@
Cognitive behavioral therapy (CBT) is a type of psychotherapeutic treatment that helps people learn how to identify and change destructive or disturbing thought patterns that negatively influence their behavior and emotions [24]. It is based on the cognitive model [25], which hypothesizes that people’s emotions, behaviors, and physiology are influenced by their perception of events. In all forms of CBT that are derived from Beck’s model [26], treatment is based on a cognitive formulation, beliefs, and behavioral strategies that characterize a specific disorder [27].
@@
CBT involves a wide range of strategies to help people overcome negative patterns, such as identifying negative thoughts, practicing new social skills, SMART goal-setting, stress problem-solving, and self-monitoring [11]. It can be used as a short-term treatment to help individuals learn to focus on present thoughts and beliefs (symptoms like addiction, anger issues, anxiety, depression, etc.) and improve mental health conditions (like chronic pain, insomnia, stress management, etc.) [24].
III Approach Overview
@@
III-A Cognitive Computing Architecture
@@ III-A1 Cloud Layer
Through receiving the analysis results from the cognitive data server, this server will learn resources about edge cloud computing, network communication, and medicine. Then, it optimizes scheduling strategies and distributes corresponding resources in real-time. Moreover, the server sends the integrated resource data back to the cognitive data server, updating its database. Specifically, the server can achieve resource optimization and energy saving by implementing the computing unload, handover strategy, caching and delivery, and the corresponding intelligent algorithms [3], which can meet various heterogeneous application requirements.
(a) Entry Level: an Example of the Assisted Robot’s Behavior Trees of “Aladdin and the Magic Lamp”.
(b) Basic Level: Assisted Robot’s Behavior Trees
(c) Middle Level: an Example of the Assisted Robot’s Behavior Trees of the Knock knock joke.
(d) Advanced Level: Examples of Assisted Robot’s Behavior Trees of the Sarcastic Jokes.
III-A2 Edge Layer
Instead of constantly moving data to the cloud for computing operations, which accounts for the energy costs, in the edge layer, data can be mined and processed on edge devices and servers closer to the user [28]. Moreover, compared with traditional edge devices, we define a novel edge intelligent robot for medical applications and smart health care, which not only can collect and process data from users and the environment but also serve as the medical equipment or tool to cure patients.
@@ III-A3 IoT Layer
III-B Cognitive Models of the Edge Intelligent Robot
@@
III-B1 Humor Styles for Cognitive Distortions
The type of humor reflects the cognitive development of people and the level of their social skills [29]. Humor styles are potential mediators of the association between cognitive and interpersonal vulnerability factors and psychological dysfunction, distress, or poor interpersonal functioning [30]. To some extent, by improving their sense of humor, ASD patients can gradually enhance their communication skills and enter the mutual socialization process with others [31].
Based on what McGhee [32] described, this research considers four stages of humor development in the whole CBT process. Using the assisted robot to design different funny scenarios and jokes treats ASD patients, especially the children, in their corresponding periods. Tab. I illustrates the four stages of human development through RAT in our research.
TABLE I: Four Stages of Humor Development through RAT
@@ III-B1 Humor Styles for Cognitive Distortions
III-B2 Behavior Tree based Humor Styles Representation
According to the above discussion, we design four specific application cases representing them as behavior trees (BT) [33] to improve patients’ experience and enhance the effect of the RAT in different stages of ASD.
@@ III-B2 Behavior Tree based Humor Styles Representation
Patients with less social skills enjoy interactions with familiar objects or games by creating something new or different out of them. Therefore, at the entry level, we design three scenarios (playing ball, chasing, and spinning) to let the assisted robot interact with patients, helping them develop interests and willingness to communicate and make new friends. Fig. 2(a) shows the three scenarios as the BT.
@@
As patients develop interests and are not resistant to communicating with new partners, they become willing to use language and gestures for fun and to engage others. There may be some overlap between Entry Level and Basic Level in humor development, but the distinguishing feature of the Basic Level is that a ”verbal statement alone creates the incongruity and leads to laughter” [32].
At this point, humor becomes a significant part of social skill development. In order to help patients gain positive emotional reactions from their partners, we create a simple pretend game [^2] Pretend game is a loosely structured form of play that generally includes role-play, object substitution, and nonliteral behavior [34]. – Aladdin and the Magic Lamp – to let them play with the robot, adding social reinforcement to the fun. Fig. 2(b) illustrates an example of the Assisted Robot’s Behavior Trees of “Aladdin and the Magic Lamp”. Here, the assisted robot plays the Magic Lamp and the patient plays the Aladdin.
@@
When patients master the above two levels of humor, their cognition will develop to Piaget’s Preoperational Stage [35]. Humor can be regarded as the intellectual play expressed through language [36], and the production and appreciation of humor also change.
In this intellectual linguistic play, the listener needs to understand the double meanings of words to find it humorous [31]. At this stage, making others laugh becomes the reward of social approval, which can drive patients to develop this social language skill. Knock-knock jokes and other ready-made jokes are popular at this level. Considering knock-knock jokes and other ready-made jokes are popular, we design the scenarios of knock-knock jokes to support patients interacting with the robot at this level (Fig. 2(c)).
@@
In the cognitive stage of concrete operations [35], patients will improve mental operations to analyze multiple aspects of a situation and perform tasks at a higher level than they could in the preoperational stage. They need to understand the meanings of puns and other forms of more abstract humor and can use inductive and deductive reasoning and reversibility in thinking about the beginning, middle, and end points of a funny story or joke [31]. Teasing and sarcastic jokes are commonly used at this level. Fig. 2(d) illustrates two examples of assisted robot’s BT of sarcastic jokes.
IV Conclusion and Future Works